(last updated on 12/21/10)
expect no beam work for most of the week. First three days are for cooling down to 4K. 2-3 days of power supply setup will follow. Due to trim quad P/S upgrade work, we can not work with beam until it is complete. There may be some EVE/OWL shifts near the end when blue is ready while P/S group is working in yellow. Estimates: ~4 shifts available for beam in blue in last two days while P/S group is working on yellow.
o AGS to RHIC synchro. The target frequency of 28.128770MHz; (RF group)
o AGS extraction and AtR tuning. (Nick Tsoupas and Vincent Schoefer)
o With Blue magnets ready (even without the hysteresis), try the injection to ensure that there is no obvious problems for circulating beam (like obstacles).
o Continue setup for Blue ring according to “Injection setup” plan in “Startup Mode” below.
Remaining P/S work continues for 3-7days during DAY and EVE shifts. Only one shift per day for beam setup..
o Set injection from ATR into RHIC rings:
§ Injection with kicker off: set bo6-tv9 and yo5-tv9 to 1.3 mrad, no snakes.
§ Injection kicker timing adjustment: (Greg Marr)
§ Minimize injection losses.
§ Steer the beam around the ring and achieve continuous circulation.
o Setup snake orbit bumps and turn on the Snakes. Adjust tunes and vertical orbits. Optimize snake bumps to eliminate injection losses at snake locations.
o RHIC RF work with 28MHz. (RF group)
o Instrumentation Setup
§ Tunemeter(Artus) (with tunes in upper half). (Angelika)
§ Injection damper setup (upper half tunes). (Rob)
§ Schottky (Kevin Brown)
§ BBQ (Kevin Mernick)
§ BTF (Kevin Mernick, Michiko)
§ IPM. (Roger Connolly, Steve)
§ WCM. (Larry Hoff)
§ CNI polarimeter: (Bill, Anatoli, Haixin)
o Catch beam with 9MHz cavity. Test the switch back and forth between 9MHz and 28MHz cavities. 1-2 shift(s).
o Machine tuning
§ Orbit correction (goal is for rms orbit below 0.3mm). (Vadim)
§
Optimize
bend trim/orbit radius.
§ Decoupling. (Yun)
§ Chromaticity correction. (Steve, Al, Michiko)
§ Optics measurement and correction. (Mei, Joe Aronson)
o Check the spin pattern request against the measured. (Bill, Anatoli, Haixin)
o Snake current scan (measuring polarization) at injection ( at least 1 ring)
o TOCC feedback development (Al, Michiko,Steve,Vadim, Yun)
o Ramp development with 6 bunches
§ Intensity: 0.15-1e11/bunch
§ Use corrector settings (skew and dipole) from last run ramp as starting point
§ Work items:
a) Radial loops.
b) tune/coupling/orbit feedback(Al, Michiko,Steve,Vadim,Yun)
c) Orbit correction along the ramp ( < 0.3mm rms )
d) Decoupling along the ramp. (Yun)
e) Chromaticity along the ramp( chromaticity ~2-3 units) (SteveT)
f) Verify transverse IR orbit separations and longitudinal relative phase. (Vadim)
o Experimental magnets, PHENIX and STAR, switch on, when ready. Do necessary ramp adjustments.
Although it is listed separately, it is expected that the 9MHz commissioning uses one and 1/2 shifts per day. The rest will be the stop development. .
o Catch beam with 9MHz cavity.
o Establish ramp with 9MHz for six bunches.
o Develop rebucketing into 28MHz.
o Step up bunch intensity (with feedback at injection).
o Step up number of bunches.
o With 6-56 bunches, bunch intensity about 1.e11
o Establishing collisions. DX BPM timing (Angelika, Rob).
o Store optics adjustments:
§ Orbit correction, decoupling, chromaticity.(Vadim, Michiko, Yun, Al)
§ Dispersion function measurements and correction (if needed) (Mei, Joe Aronson)
§ IR nonlinear corrections. (Chris Zimmer)
§ Betatron tune scans: verifying available working point window.
§ Optics measurements with AC dipole. (Mei, Joe Aronson)
o Collimation setup. (Angelika)
o Evaluation of store reproducibility and slow orbit feedback. Optics comparisons for stores separated by 24h. (Vadim, Michiko)
o Evaluation of beam-beam limit:
§ (Gradual) increase in bunch intensity (goal is 1.4e11)
§ Observing lifetime, luminosity and experiment background dependence on the bunch intensity.
§ Measuring beam-beam tune shifts and emittance blowups.
o Good polarization transmission efficiency on the ramp.
§ Tune and orbit adjustments on the ramp.
§ Polarization measurements on the ramp.(Bill, Anatoli, Haixin)
§ Scan vertical tune on the ramp with different rms orbit. Check polarization on the ramp and at store.
o Testing 10Hz feedback system (depends on preparedness) (Rob, Michiko)
o Test if the 197MHz cavities can be used to reduce the vertex further. (RF group)
o Rotator ramp development. (Waldo)
2 shifts per day; Owl shift for experiment setup.
o 109 bunches. Verify injection kicker timing. Ensure that there is no emittance blowup. (Greg)
o Evaluate pressure rise at the injection with 109 bunches and more than 1.3e11/bunch. (Wolfram, SY)
o Do beam conditioning of vacuum pipe with intense beam, if needed (Wolfram,SY)
o 109 bunch ramp development:
§ (Gradually) increase intensity bu increase bunch numbers, starting with ~1.4e11 per bunch.
§ Work on beam losses reduction (orbit, tunes adjustments) on the ramp. Evaluating the intensity limit due to beam losses (if exists)
§ Verify relative longitudinal phase.
§ Collimation on the ramp (in some form). (Angelika)
o CNI polarimeter tuning. (Bill, Anatoli, Haixin)
o Detector background evaluation. (Angelika)
o Make judgments about optimal operation parameters: number of bunches, bunch intensities and beta*
o Optimization of the beginning-of-store activities. (Angelika)
o First week: machine development on DAY shift if needed
o After first week: ~ 8 hours machine development on Thursday, to fix identified urgent machine problems and commissioing new system.